clear fractions and decimals, solve and check
1/2+4m=3m-5/2
i am really stuck on how to solve it properly the denominater 2 is under the 1 and the denominater 2 is under the 5

Found 3 solutions by stanbon, mikerueckert, Nate:
Answer by stanbon(75887)

1/2+4m=3m-5/2
Is the denominator on the left side "2" or "2+4m"?
Is the numerator on the right side "5" or "3m-5"?
Using parentheses properly can clarify these postings.
Cheers,
Stan H.

Answer by mikerueckert(1)
Answer by Nate(3500)

1/2 + 4m = 3m - 5/2
1/2 + 4m - 3m = 3m - 3m - 5/2
1/2 + m = -5/2
1/2 - 1/2 + m = -5/2 - 1/2
m = -6/2
m = -3